OpenStudy (anonymous):

Solve using perfect square factoring patterns. x2 – 10x + 25 = 9 A. {4, –12} B. {2, 8} C. {–8, –2} D. {–4, 12}

OpenStudy (sparrow2):

(x-5)^2=9 x-5=3 x=8 or x-5=-3 x=2